OpenStudy (anonymous):

Arithmetic sequence. sequence 1: 2, 6,...... sequence 2: 24, 12,..... sequence 3: 1, 5,...... Assuming that the sequences above are arithmetic with t(1) as the first term, find the next four terms for each sequence. For each sequence, write an explanation of what you did to get the next term and write an equation for t(n).

OpenStudy (anonymous):

@jim_thompson5910 @dan815 @Nnesha

jimthompson5910 (jim_thompson5910):

sequence 1: first term is 2, second is 6 since 6-2 = 4, we are adding 4 to each term to get the next term so the third term is 6+4 = 10 the fourth term is 10+4 = 14 etc etc

OpenStudy (anonymous):

how would you put in an equation tho @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you use t(n) = t(1) + d(n-1) that's the general arithmetic formula

jimthompson5910 (jim_thompson5910):

for sequence 1, t(1) = 2 is the first term d = 4 is the common difference

OpenStudy (anonymous):

what do i put for n?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do you get for t(n) ?

OpenStudy (anonymous):

well what do i put for d(n-1), what do i put for this n? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you leave n as it is

jimthompson5910 (jim_thompson5910):

you'll get an expression in terms of n

OpenStudy (anonymous):

you get t(n) = 4n-1

jimthompson5910 (jim_thompson5910):

what is t(n) equal to when n = 1?

OpenStudy (anonymous):

2?

jimthompson5910 (jim_thompson5910):

t(n) = 4n-1 t(1) = 4(1)-1 t(1) = ??

OpenStudy (anonymous):

t(1) + d(n-1) 2 + 4(1-1) 2 + 0 = 2 i thought you did this @jim_thompson5910

jimthompson5910 (jim_thompson5910):

it works there, but it doesn't work with t(n) = 4n-1

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i get it thnx cx